Grace Co. can further process Product B to produce Product C. Product B is currently selling for $60 per pound and costs $38 per pound to produce. Product C would sell for $95 per pound and would require an additional cost of $13 per pound to produce. What is the differential revenue of producing and selling Product C?
Answer #1
Concepts and reason

Differential revenue: It refers to the distinction in sales that will be created by two different courses of action. The concept of differential revenue is commonly used to evaluate which of the two or more investments to be invested in a business.

Fundamentals

Sales: Sale of any goods or services can be made on a cash or credit basis. The amount receivable on sale can either be received immediately in cash or such a payment can be received at some future date. In case of sale is being made on a credit basis the company maintains an account of such customer in its books as Debtor or Accounts Receivable.

Product Cost: It is the cost which is associated with the production of goods. If there is no production then there would be no product cost. The examples of product costs are direct material, direct labor, and factory overhead.

Differential cost: It is the difference in the total cost between the two mutually exclusive options for consideration. This method is also used in situations where there are step costs involved where the production of one more unit will result in additional cost.

Cost and Expenses: Cost is the amount expended to any particular product or amount related to any particular product. Whereas, expenses means the cost incurred in order to earn income. Prepaid Expense and Outstanding Expense are its two types. It is shown on the debit side of profit and loss Account

The concept of differential revenue is used in determining the best utilization of the resources of the company.

Compute the differential revenue of producing and selling product C using the equation as shown below:

Differentialrevenue=(RevenuefromsaleofproductCRevenuefromsaleofproductB)=$95$60=$35\begin{array}{c}\\{\rm{Differential revenue}} = \left( \begin{array}{l}\\{\rm{Revenue from sale of product C}} - \\\\{\rm{Revenue from sale of product B}}\\\end{array} \right)\\\\ = \$ 95 - \$ 60\\\\ = \$ 35\\\end{array}

Ans:

The differential revenue of producing and selling of product C is $35.